Part A If the voltage across a circuit of constant resistance is doubled, how is the...

Part A

If the voltage across a circuit of constant resistance is doubled, how is the current in the circuit affected? If the voltage across a circuit of constant resistance is doubled, how is the current in the circuit affected?

1-The current is reduced by a factor of 4.

2-The current is doubled.

3- The current is reduced by a factor of 2.

4-The current remains constant.

5-The current is quadrupled.
Answer #1
Concept and reason

The concept required to solve this question is Ohm’s law.

First, write the expression for the new current in the circuit by using the Ohm’s law. Finally, determine the variation of the current in the circuit with voltage.

Fundamentals

According to ohm’s law, the current in the circuit is directly proportional to the voltage across a circuit.

ViV \propto i

The expression for the voltage across the circuit is,

V=iRV = iR

Here, VV is the voltage across the circuit, ii is current in the circuit, and RR is the resistance of the material.

(A)

The relation between voltage and current is,

V=iRV = iR

The relation between the initial voltage V1{V_1} across the circuit and the initial current i1{i_1} in the circuit when the resistance remains constant is,

V1=i1R{V_1} = {i_1}R

The relation between the final voltage V2{V_2} across the circuit and the final current i2{i_2} in the circuit when the resistance remains constant is,

V2=i2R{V_2} = {i_2}R

Solve the above two equations for the new current i2{i_2} in the circuit.

V2V1=i2Ri1RV2V1=i2i1i2=(V2V1)i1\begin{array}{c}\\\frac{{{V_2}}}{{{V_1}}} = \frac{{{i_2}R}}{{{i_1}R}}\\\\\frac{{{V_2}}}{{{V_1}}} = \frac{{{i_2}}}{{{i_1}}}\\\\{i_2} = \left( {\frac{{{V_2}}}{{{V_1}}}} \right){i_1}\\\end{array}

The expression for the current i2{i_2} in terms of current i1{i_1} is,

i2=(V2V1)i1{i_2} = \left( {\frac{{{V_2}}}{{{V_1}}}} \right){i_1}

Given that, the voltage across the circuit is doubled, that is V2=2V1{V_2} = 2{V_1} .

Substitute 2V12{V_1} for V2{V_2} in equation i2=(V2V1)i1{i_2} = \left( {\frac{{{V_2}}}{{{V_1}}}} \right){i_1} .

i2=(2V1V1)i1=2i1\begin{array}{c}\\{i_2} = \left( {\frac{{2{V_1}}}{{{V_1}}}} \right){i_1}\\\\ = 2{i_1}\\\end{array}

Ans: Part A

The current is doubled when the voltage across the circuit is doubled.